# Simulation-trained conditional normalizing flows for likelihood approximation: a case study in stress regulation kinetics in yeast

### v1

@@ Abstract
Physics-inspired inference often hinges on the ability to construct a likelihood, or the probability of observing a sequence of data given a model. These likelihoods can be directly maximized for parameter estimation, incorporated into Bayesian frameworks, or even used as loss functions in neural networks. Yet, many models, despite being conceptually simple, lack tractable likelihoods. A notable example arises in estimating protein production from snapshot measurements of actively dividing cells. Here, the challenge stems from cell divisions occurring at non-Exponentially distributed intervals with each division stochastically partitioning protein content between daughter cells, making protein counts in any given cell a function of its full division history. Such history dependence precludes a straightforward likelihood based on a (standard Markovian) master equation. Instead, we employ conditional normalizing flows (a class of neural network models designed to learn probability distributions) to approximate otherwise intractable likelihoods from simulated data. As a case study, we examine activation of the glc3 gene in yeast involved in glycogen synthesis and expressed under nutrient-limiting conditions. We monitor this activity using snapshot fluorescence measurements via flow cytometry, where GFP expression reflects glc3 promoter activity. A naïve analysis of flow cytometry data ignoring cell division suggests many cells are active with low expression. However, fluorescent proteins persist and can be inherited, so cells may appear active from retaining ancestral fluorescence. Explicitly accounting for the (non-Markovian) effects of cell division reveals glc3 is mostly inactive under stress, showing that while cells occasionally activate it, expression is brief and transient.
@@ 1 Introduction
In the natural sciences, we learn to simulate dynamical models, whether deterministic or stochastic, in order to replicate experimental output and eventually predict phenomena. These models typically depend on parameters (e.g., forces or velocity fields) that can be tuned to approximate the experimental output we wish to simulate. Simulations like these have provided significant insights across scientific areas: from reactions in solution [1, 2], to gene expression [3, 4], neuronal activity [5, 6], dynamics of infection [7, 8], and even toward explaining processes involved in scientific collaboration and publishing [9, 10]. Learning parameters from data, usually termed “inverse modeling”, however, demands more than the ability to simulate data from given parameters. To rigorously learn parameters, one typically formulates likelihoods — probability densities for observing data sequences given parameter sets ($\theta$), $p(\text{data}|\theta)$. Using such likelihoods, parameters are then estimated: either by finding those parameters making the sequence of the data’s likelihood largest (from which derives the notion of maximum likelihood); or by constructing posteriors over the parameters informed by the data from the likelihood [11, 12, 13, 14, 9].
@@
Consider, for sake of example, an even more challenging case: we monitor cells at various points in time (snapshot data). From this snapshot data, our goal is to estimate the production of a particular (labeled) protein species through fluorescence. It is common to model the dynamics of this protein’s number using Markovian dynamics [19, 20, 21, 22, 23, 24, 25, 26]. For this iteration of the problem, inferring parameters is a textbook problem [13]. However, when experiments span multiple cell division cycles, the Markovian assumption fails, as many proteins have half-lives longer than a single division [27, 28, 29]. As such, reduction in protein numbers occurs, mainly, via cell division rather than degradation. Since, cell division times are not exponentially distributed [30, 31], the cell cycle cannot be modeled using Markovian dynamics. While it is possible to simulate protein levels under cell division [32, 33, 34], to our knowledge, no established inference framework allows us to write down a general likelihood for this problem.
One idea to bypass likelihood calculation is to use (forward) simulations to approximate the likelihood, as the simulated results can be understood as samples from $p(\text{data}|\theta)$, leading to what is known as simulation-based inference (SBI) [35, 36, 37]. An early implementation of SBI is approximate Bayesian computation (ABC), where the likelihood is approximated by the fraction of simulated data points falling within a threshold distance, $\epsilon$, of observed data [38]. However, ABC presents limitations: (1) it requires new simulations for each parameter values $\theta$, making likelihood evaluation computationally intensive; (2) it depends on arbitrary distance metrics; and (3) likelihood estimates can vary unpredictably, particularly for rare events or large datasets.
Recent advances in neural networks have inspired new approaches for SBI. Some of these consist of finding a neural network approximation of the Bayesian posterior directly, bypassing the likelihood altogether [39, 40, 41, 42]. However, without an explicit likelihood function, we lose a key tool in validating the inferred posterior. When a likelihood is available, we can compare the approximated likelihood with histograms obtained from simulations in regions of high posterior. In some particular benchmarks [43, 44], it has been argued that posterior approximations can produce overconfident posteriors underestimating uncertainty. Without an estimated likelihood to compare against, problems like these can go unnoticed. As a result, purely posterior-focused approaches have faced criticism for limited generalizability and the lack of reliable validation strategies [45, 43, 44].
@@
While conditional flows have been proposed for general density estimation tasks [42, 54], their application as intractable likelihood approximators remains relatively underexplored in scientific modeling with recent work concentrated on classification problems [55]. In this work, we broaden their scope by applying them to more general and interpretable scientific models and, in doing so, uncover the kinetics of state switching in yeast (S. cerevisiae) by accurately taking into account proteins counts across cell division.
@@
To demonstrate the flexibility of our approach, we apply it to a hierarchy of three increasingly realistic biological models. We begin with a simple system in which a cell produces protein of interest at a constant rate and divides deterministically after fixed time intervals which we will later define as Model 1. Next, in Model 2, we extend Model 1 to include stochastic cell division times. Finally, in Model 3 we consider a realistic experimental setting where levels of activation of the glc3 gene in yeast, involved in glycogen synthesis, are monitored under nutrient-limiting conditions via fluorescent flow cytometry. By considering cell division in our inference on glc3 activity, we uncover that even under high stress conditions, glc3 remains inactive in nearly all cells with rapid switches to an upregulated state. For concreteness, a high level SBI workflow and description of our models are summarized in Fig. 1.
@@ 2.2 Normalizing flows for likelihood approximation
As a result, minimizing the KL divergence is equivalent to minimizing the empirical negative log-likelihood of the simulated data
@@ 4 Conclusion
Inference is a critical step in going from data to a parametrized model. Yet physics-inspired inference is limited by our ability to mathematically formulate likelihoods, the probability of observing the data given the model. This often reduces to solving the coinciding master equation [17, 3, 4, 16], Fokker-Planck [11, 82, 83] and other often equations which often cannot be solved or easily formulated though the process can be simulated.
In this work, we demonstrated how neural networks can be leveraged as approximators to likelihoods using data from forward simulations in inference problems relevant not only to simple to simulate yet pathologically non-Markovian dynamics but also to inference problems where the likelihood is computationally expensive. In such settings, the method presented here provides a practical surrogate preserving the statistical framework of likelihood-based inference while substantially reducing computational costs.
While we have opted for normalizing flows in this work due to their straightforward probabilistic formulation, though other neural network architectures whose outputs are tractable probability distributions (such as score-based diffusion models [48]) may also have been used. For example, diffusion models have been applied to infer Boltzmann weights for proteins [84], offering a promising alternative generative modeling framework.
As an illustrative example, we applied this approach to a non-Markovian biological system involving protein production kinetics, where protein abundance is reduced through cell division, thereby making the instantaneous protein count within a cell dependent on its complex cell-division history.
This feature essentially undermines simplifying assumptions used thus far to make this inference problem tractable such as treating cell division on protein abundance is frequently modeled as an additional chemical reaction [19, 20, 21, 22, 23, 24, 25, 26] which assumes that reduction in protein numbers arises solely from individual degradation rather than partitioning during inheritance, in a non-Markovian cell cycle. As shown in Fig. 2, such approximations can lead to qualitatively incorrect likelihood distributions. In contrast, our method models cell division times explicitly and performs likelihood-based inference directly from forward simulations, thereby avoiding these oversimplifications.
We applied this likelihood approximation technique to both deterministic (Model 1) and stochastic (Model 2) models of cell division time, as shown in Figs. 3 and 4, respectively. In the deterministic case, the method accurately captured the steady-state protein distribution described in (15), demonstrating its ability to learn complex and otherwise intractable likelihood functions. When extended to stochastic division times, the method also proved effective in multi-parameter inference, successfully estimating both the protein production rate, $\beta$, and the division time variability, $\sigma$.
We further incorporated additional biological complexity, including two-state gene regulation and asymmetric cell division, generalizing inference to more realistic scenarios such as those found in flow cytometry datasets of S. cerevisiae. The extended model allowed us to infer phenotypic gene activation and inactivation rates, $\lambda_{\text{act}}$ and $\lambda_{\text{ina}}$, from synthetic data (Fig. 5).
@@
We then applied this framework to analyze real flow cytometry data from S. cerevisiae cultured in chemostats at different dilution and thus nutrient concentration rates (Fig. 6). The results show that contrary to expectation — and although fluorescence intensity is significantly higher under high-stress conditions — cells rarely find themselves in a glc3 upregulated state [57, 76] instead leveraging gene product inherited through parent cells. Interpreting these results without accounting for inheritance across the cell cycle would lead to fundamentally different conclusions, suggesting that glc3 is frequently upregulated ignoring altogether gene products inherited through division.
@@ Acknowledgments
MCMC initialization in Section 3.4.
For the MCMC results presented in Section 3.4, we begin by applying an approximate Bayesian computation (ABC) procedure. In this phase, we simulate data under the model for each proposed parameter value and compute a distance to the observed data based on intensity histograms. A procedure similar to the one described above is then applied to generate 10,000 accepted samples. This initial sampling takes approximately four hours but is not reliable: the likelihood estimates for a fixed parameter value often vary by orders of magnitude, and the method tends to identify only a restricted region of parameter space where the observed fluorescence intensities are approximately matched. Nonetheless, we use the mean and standard deviation of these samples to define a Gaussian prior for the subsequent MCMC sampling described above.
@@
This section describes the data preprocessing pipeline applied to the flow cytometry measurements discussed in Section 3.4 and visualized in Figure 6a. The goal of this cleaning procedure is to isolate a consistent population of cells suitable for downstream analysis by removing artifacts and outliers that arise during acquisition or due to biological heterogeneity.
@@
While more complex models (e.g., multiple or non-parametric states) are conceivable, they must be reflected in the observable fluorescence data. Here we argue that it is reasonable to assume no production in the inactive state. In the case where there is a single producing state, as described in Sec. 3.1, the expected steady-state protein number $G$ in Eq. (15) is given by $\langle G\rangle=\frac{3}{2}{\beta}{T}$, where $T$ is the cell division time. In chemostats, where division is set by the dilution rate $D$ (see SI S1), this becomes $\langle G\rangle=\frac{3}{2}\frac{\beta}{D}$. If cells were always producing, slower dilution (and thus longer division times) would lead to higher average protein numbers and a shift in the fluorescence peak.
@@
In Section 3.4, we mention that training the neural network on unrepresentative regions of parameter space can bias inference. To mitigate this, we initialize training using a prior informed by a simpler SBI method, Approximate Bayesian Computation (ABC). This method estimates a rough likelihood by comparing histograms of simulated flow cytometry intensity values at various parameters $\theta$ to the observed data shown in Fig. 6. We then perform the MCMC sampling procedure detailed in Sec. S4. We use this MCMC sampled values to construct the prior for Sec. 3.4. The specifics of the ABC approximation are provided in Sec. S7.1, while Sec. S7.2 explains why our normalizing flows-based approach yields a more reliable posterior.

### v2

# Inherited or produced? Inferring protein production kinetics when protein counts are shaped by a cell’s division history

@@ Abstract
Inferring protein production kinetics for dividing cells is complicated protein inheritance from the mother cell. For instance, fluorescence measurements–commonly used to assess gene activation–may reflect not only newly produced proteins but also those inherited through successive cell divisions. In such cases, observed protein levels in any given cell are shaped by its division history. As a case study, we examine activation of the glc3 gene in yeast involved in glycogen synthesis and expressed under nutrient-limiting conditions. We monitor this activity using snapshot fluorescence measurements via flow cytometry, where GFP expression reflects glc3 promoter activity. A naïve analysis of flow cytometry data ignoring cell division suggests many cells are active with low expression. Explicitly accounting for the (non-Markovian) effects of cell division and protein inheritance makes it impossible to write down a tractable likelihood–a key ingredient in physics-inspired inference, defining the probability of observing data given a model. The dependence on a cell’s division history breaks the assumptions of standard (Markovian) master equations, rendering traditional likelihood-based approaches inapplicable. Instead, we adapt conditional normalizing flows (a class of neural network models designed to learn probability distributions) to approximate otherwise intractable likelihoods from simulated data. In doing so, we find that glc3 is mostly inactive under stress, showing that while cells occasionally activate the gene, expression is brief and transient.
Keywords: Transcriptional kinetics, Non-Markovian dynamical systems, Bayesian inference, Proteomics, Cell cycle, Neural Networks.
@@ 1 Introduction
Here we envision an experiment monitoring cells at various points in time (snapshot data). From this snapshot data, our goal is to estimate the production of a particular (labeled) protein species through fluorescence. It is common to model the dynamics of this protein’s number using memoryless (Markovian) dynamics [1, 2, 3, 4, 5, 6, 7, 8]. For this iteration of the problem, inferring parameters such as protein production or degradation rates is straughforward and relies on our ability to formulate a likelihood [9]. Indeed, learning parameters from data, usually termed “inverse modeling”, generically requires one to formulate likelihoods — probability densities for observing data sequences given parameter sets ($\theta$), $p(\text{data}|\theta)$. Using such likelihoods, parameters are then estimated: either by finding those parameters making the sequence of the likelihood largest (from which derives the notion of maximum likelihood); or by constructing posteriors over the parameters using the likelihood [10, 11, 9, 12, 13].
However, when experiments span multiple cell division cycles, a fundamental modeling challenge occurs as many proteins have half-lives longer than a single division [14, 15, 16]. As such, reduction in protein numbers occurs, mainly, via cell division rather than degradation. Protein counts in any given cell thus depend on the cell’s division history. Put differently, the Markovian assumption–inherent to existing modeling strategies–fails as, to our knowledge, no established inference framework allows us to write down likelihoods in such a general setting. Since, cell division times are not exponentially distributed [17, 18], the cell cycle cannot be modeled using Markovian dynamics. Yet, despite the challenges in writing down a likelihood, it is possible to simulate protein levels under cell division [19, 20, 21].
For this reason, we propose an alternative paradigm to infer kinetics of protein production in the presence of cell division: the ability to perform inference assuming we can simulate processes but cannot write down the likelihood [22, 23, 24].
Indeed, in the natural sciences, we learn to simulate dynamical models, whether deterministic or stochastic, in order to replicate experimental output and eventually predict phenomena. These models typically depend on parameters (e.g., forces or velocity fields) that can be tuned to approximate the experimental output we wish to simulate. Simulations like these have provided significant insights across scientific areas: from reactions in solution [25, 26], to gene expression [27, 28], neuronal activity [29, 30], dynamics of infection [31, 32], and even toward explaining processes involved in scientific collaboration and publishing [13, 33].
@@
For the case where they cannot even be formulated, one idea to bypass likelihood calculation using (forward) simulations to approximate the likelihood, as the simulated results can be understood as samples from $p(\text{data}|\theta)$. This leads to what is known as simulation-based inference (SBI) [23, 22, 24]. An early implementation of SBI is approximate Bayesian computation (ABC), where the likelihood is approximated by the fraction of simulated data points falling within a threshold distance, $\epsilon$, of observed data [38]. However, ABC presents a number of limitation true to some ABC implementations: (1) it requires new simulations for each parameter values $\theta$, making likelihood evaluation computationally intensive; (2) it depends on arbitrary distance metrics; and (3) likelihood estimates can vary unpredictably, particularly for rare events or large datasets.
To avoid these limitations, recent advances in neural networks have inspired new approaches for SBI. Some of these consist of finding a neural network approximation of the Bayesian posterior directly, bypassing the likelihood altogether [39, 40, 41, 42]. However, without an explicit likelihood function, we lose a key tool in validating the inferred posterior. When a likelihood is available, we can compare the approximated likelihood with histograms obtained from simulations in regions of high posterior. In some particular benchmarks [43, 44], it has been argued that posterior approximations can produce overconfident posteriors underestimating uncertainty. Without an estimated likelihood to compare against, problems like these can go unnoticed. As a result, purely posterior-focused approaches have faced criticism for limited generalizability and the lack of reliable validation strategies [45, 43, 44].
@@
While conditional flows have been proposed for general density estimation tasks [42, 54], their application as intractable likelihood approximators remains relatively underexplored in scientific modeling with recent work concentrated on classification problems [55]. In this work, we broaden their scope by applying them to more general and interpretable scientific models and, in doing so, uncover the kinetics of gene state switching in yeast (S. cerevisiae) by accurately taking into account proteins counts across cell division.
@@
To demonstrate the flexibility of our approach and its eventual applicability to inference of protein production kinetics under cell division, we apply it to a hierarchy of three increasingly realistic biological models. We begin with a simple system in which a cell produces protein of interest at a constant rate and divides deterministically after fixed time intervals which we will later define as Model 1. Next, in Model 2, we extend Model 1 to include stochastic cell division times. Finally, in Model 3 we consider a realistic experimental setting where levels of activation of the glc3 gene in yeast, involved in glycogen synthesis, are monitored under nutrient-limiting conditions via fluorescent flow cytometry. By considering cell division in our inference on glc3 activity, we uncover contrary to expectation that even under high stress conditions, glc3 remains inactive in nearly all cells with rapid switches to an upregulated state. For concreteness, a high level SBI workflow and description of our models are summarized in Fig. 1.
@@ 2.2 Normalizing flows for likelihood approximation
As a result, minimizing the KL divergence in (7) is equivalent to minimizing the empirical negative log-likelihood of the simulated data
@@ 4 Conclusion
We considered the challenge of estimating protein production, even extending to the case of gene switching kinetics, when cells are dividing. In this case, the instantaneous protein count within each cell depends on its entire cell-division history.
@@
We applied this framework to real flow cytometry data from S. cerevisiae cultured in chemostats at different dilution and thus nutrient concentration rates (Fig. 6). The results show that contrary to expectation — and although fluorescence intensity is significantly higher under high-stress conditions — cells rarely find themselves in a glc3 upregulated state [57, 76] leveraging instead gene product inheritance through parent cells. Interpreting these results while ignoring the cell cycle would lead to fundamentally different conclusions: glc3 is frequently upregulated ignoring altogether gene products inherited through division.
Indeed, our treatment now extends approaches that so far have made the simplifying assumption of modeling cell division as an additional chemical reaction [1, 2, 3, 4, 5, 6, 7, 8] implicitly assuming that reduction in protein numbers arises solely from individual degradation rather than partitioning. As shown in Fig. 2, such approximations can lead to qualitatively incorrect likelihood distributions. In contrast, our method models cell division times explicitly and performs likelihood-based inference directly from forward simulations, thereby avoiding these oversimplifications.
Learning protein production and gene switching kinetics requires we move beyond existing physics-inspired inference schemes relying on likelihoods which, in turn, often reduces to solving the coinciding master equation [36, 27, 28, 35], Fokker-Planck [10, 82, 83] and other often equations which often cannot be solved or easily formulated though the process may be simulated.
In this work, we motivated how neural networks can be leveraged as approximators to likelihoods using data from forward simulations for both common (yet intractable non-Markovian) problems but, potentially, also for problems where the likelihood is computationally expensive without necessarily being fundamentally intractable. In such settings, the method presented here provides a practical surrogate preserving the statistical framework of likelihood-based inference while substantially reducing computational costs.
While we have opted for normalizing flows in this work due to their straightforward probabilistic formulation, other neural network architectures whose outputs are tractable probability distributions (such as score-based diffusion models [48]) may also have been used. For example, diffusion models have been applied to infer Boltzmann weights for proteins [84], offering a promising alternative generative modeling framework.
@@ Acknowledgments
MCMC initialization in Sec.  3.4.
For the MCMC results presented in Sec.  3.4, we begin by applying an approximate Bayesian computation (ABC) procedure. In this phase, we simulate data under the model for each proposed parameter value and compute a distance to the observed data based on intensity histograms. A procedure similar to the one described above is then applied to generate 10,000 accepted samples. This initial sampling takes approximately four hours but is not reliable: the likelihood estimates for a fixed parameter value often vary by orders of magnitude, and the method tends to identify only a restricted region of parameter space where the observed fluorescence intensities are approximately matched. Nonetheless, we use the mean and standard deviation of these samples to define a Gaussian prior for the subsequent MCMC sampling described above.
@@
This section describes the data preprocessing pipeline applied to the flow cytometry measurements discussed in Sec.  3.4 and visualized in Figure 6a. The goal of this cleaning procedure is to isolate a consistent population of cells suitable for downstream analysis by removing artifacts and outliers that arise during acquisition or due to biological heterogeneity.
@@
While more complex models (e.g., multiple or non-parametric states) are conceivable, they must be reflected in the observable fluorescence data. Here we argue that it is reasonable to assume no production in the inactive state. In the case where there is a single producing state, as described in Sec. 3.1, the expected steady-state protein number $G$ in (15) is given by $\langle G\rangle=\frac{3}{2}{\beta}{T}$, where $T$ is the cell division time. In chemostats, where division is set by the dilution rate $D$ (see SI S1), this becomes $\langle G\rangle=\frac{3}{2}\frac{\beta}{D}$. If cells were always producing, slower dilution (and thus longer division times) would lead to higher average protein numbers and a shift in the fluorescence peak.
@@
In Sec. 3.4, we mention that training the neural network on unrepresentative regions of parameter space can bias inference. To mitigate this, we initialize training using a prior informed by a simpler SBI method, Approximate Bayesian Computation (ABC). This method estimates a rough likelihood by comparing histograms of simulated flow cytometry intensity values at various parameters $\theta$ to the observed data shown in Fig. 6. We then perform the MCMC sampling procedure detailed in Sec. S4. We use this MCMC sampled values to construct the prior for Sec. 3.4. The specifics of the ABC approximation are provided in Sec. S7.1, while Sec. S7.2 explains why our normalizing flows-based approach yields a more reliable posterior.